OpenStudy (mathisfun13):

For the geometric series 1+4+16+64+256 what is the value of r?

OpenStudy (mathisfun13):

i think its 341

jimthompson5910 (jim_thompson5910):

To find r, just pick any term (but the first one) and divide it by the previous term

ganeshie8 (ganeshie8):

its not asking for sum @mathisfun13

ganeshie8 (ganeshie8):

its asking just for the common ratio r

OpenStudy (mathisfun13):

256?

ganeshie8 (ganeshie8):

nopes. do exactly as jim suggested

ganeshie8 (ganeshie8):

if u take 256, its previous term is 64. so r = 256/64 = ?

OpenStudy (mathisfun13):

4

ganeshie8 (ganeshie8):

yes, if you take 16, its previous term is 4 so r = 16/4 = ?

OpenStudy (mathisfun13):

4

OpenStudy (mathisfun13):

so the value of r is 4

ganeshie8 (ganeshie8):

yes you can try that wid any term and its previous term. you wil end up wid 4 always

ganeshie8 (ganeshie8):

yes ! r is the common ratio

ganeshie8 (ganeshie8):

equals 4 here